# Dynamic spillovers and investment strategies across artificial intelligence ETFs, artificial intelligence tokens, and green markets

### v1

@@ Abstract
This paper investigates the risk spillovers among AI ETFs, AI tokens, and green markets using the $R^{2}$ decomposition method. We reveal several key insights. First, the overall transmission connectedness index (TCI) closely aligns with the contemporaneous TCI, while the lagged TCI is significantly lower. Second, AI ETFs and clean energy act as risk transmitters, whereas AI tokens and green bond function as risk receivers. Third, AI tokens are difficult to hedge and provide limited hedging ability compared to AI ETFs and green assets. However, multivariate portfolios effectively reduce AI tokens investment risk. Among them, the minimum correlation portfolio outperforms the minimum variance and minimum connectedness portfolios.
@@ 1 Introduction
The onset of the Fourth Industrial Revolution has ushered human society into the era of artificial intelligence (AI). This transformation has not only revolutionized daily life, work practices, and communication methods but has also profoundly impacted education, healthcare, finance, and transportation [1]. Simultaneously, technological advancements have facilitated the evolution of financial markets, leading to the emergence of artificial intelligence related financial instruments. AI-related assets refer to a category of financial instruments closely associated with AI technology and its industrial ecosystem, encompassing AI-themed exchange-traded funds (AI ETFs), AI-related publicly traded stocks, and AI-integrated cryptocurrency assets (AI Tokens). AI ETFs allocate capital to companies in AI technology, robotics, cloud computing, and automation, offering diversified exposure to the AI sector [2]. AI-related stocks include companies where AI is a core business component or holds a significant role in revenue and strategic positioning, such as NVIDIA and Google [3, 4]. And AI Tokens integrate AI and blockchain technology, primarily facilitating decentralized AI computing, data sharing, and smart contract execution to support distributed AI development [3, 2]. These innovative asset classes not only represent the forefront of technological progress but have also become novel investment instruments, attracting significant interest from institutional and individual investors [2].
The interconnections among AI-related assets and and other instruments have also garnered significant attention, including stocks, bonds, commodities, cryptocurrencies, and electricity [2, 5, 4, 6, 7, 8, 9, 10, 11, 12]. Under normal market conditions, AI tokens can serve as effective diversifiers within portfolios containing traditional assets. However, during periods of extreme market stress, neither AI tokens nor AI ETFs provide significant risk diversification benefits for other asset classes. Moreover, AI tokens have the potential to provide cost-effective hedging for traditional assets [2]. The connectedness between AI assets and fossil fuel markets varies across time-frequencies and quantiles, with short-term fluctuations driving shock transmission, while long-term dynamics can shift assets between net transmitter and receiver roles [3]. The directional predictability between bitcoin, fintech, and AI stocks fluctuates across different time lags [6].
A growing body of research explores the interdependence among different types of AI-based assets [2, 9, 3, 13]. Directional effects indicate fluctuating connectedness across market indices, with AI stocks as persistent net transmitters, and AI tokens as primary shock recipients [9]. However, the correlation between AI stocks and tokens tends to be relatively low and may vary over time [13]. AI ETFs consistently act as net transmitters, dominating the system across market conditions, while AI tokens serve as weak net recipients at median quantiles, with all variables alternating between net transmitters and recipients over time [2]. SingularityNET (AGIX) and Fetch.ai (FET), two AI related tokens, consistently act as net recipients of volatility spillovers across all three frequency domains, while other markets assume varying roles depending on the investment horizon [3].
Some scholars have examined the determinants of spillovers and interactions between AI-related assets and other asset classes. Generally, the impact of the factors on spillover effects varies across different market conditions [4]. At the 5th and 95th percentiles, financial, economic, and uncertainty variables exhibit a highly similar impact on the spillover effects among AI-based assets, sukuk, and islamic equity indices [9]. Economic policy uncertainty and U.S. stock market volatility increase the interconnectedness between AI and fintech stocks, while geopolitical risk and gold market fluctuations have a negative impact on their correlations [11].
However, the rapid growth of AI technology, driven by deep learning and cloud computing, has sharply increased global energy demand. Its reliance on fossil fuels makes it a contributor to carbon emissions and a source of environmental concern [14]. Amid efforts to pursue a low-carbon economy, financial markets are increasingly integrating environmental factors into investment decisions and financing green projects, thus spurring the rise of green assets. Meanwhile, green assets also offer investors diversification and risk management opportunities [15]. Moreover, there exits potential connections between AI- related assets and sustainable markets. The empirical findings indicate that the AI index consistently emits shocks to clean markets across all frequencies, while the S&P Global Clean Energy Index functions as a net sender of volatility spillover across all market conditions and frequencies [7]. Investors incorporating stocks of artificial intelligence and robotics companies, cryptocurrencies, and green bonds into their portfolios should be aware that certain investment risks persist [10].
Although existing research has extensively explored the correlations between AI-related assets and other financial assets, studies on their interdependence with green markets remain relatively scarce. However, the development of green markets plays a crucial role in advancing the low-carbon economic transition and improving asset allocation efficiency. Furthermore, limited research has explored the role of sustainability-linked financial instruments in hedging risks and optimizing portfolio performance in relation to AI assets. Moreover, AI tokens, as a unique type of crypto asset, have introduced significant changes in global financial dynamics. However, research on their broader impact remains limited.
This study aims to quantify risk spillovers among AI ETFs, AI tokens, and green assets via the $R^{2}$ decompostion method [16, 17]. Further, we estimate the bilateral hedging ratios and hedge effectiveness of different assets pairs. Additionally, we apply the the minimum variance portfolio (MVP) [18], and compare its performance with the minimum correlation portfolio (MCP) [19], and the minimum connectedness portfolio (MCoP) [20]. This paper has the following contributions: First, examining risk spillovers among AI tokens, AI ETFs, and green assets helps uncover cross-market risk transmission patterns and potential contagion effects. This paper expands the existing literature by providing new insights into their interdependence. Second, by distinguishing between contemporaneous and lagged risk spillovers, this study provides a more comprehensive understanding of asset interdependencies over time. Third, the findings offer valuable insights for investors in optimizing asset allocation. Moreover, this study assists regulators in assessing market stability with a structured framework for risk transmission dynamics among AI and green assets.
@@
The remainder of this paper is structured as follows. Section 2 provides a description of the data sets, along with summary statistics. Section 3 provides the methodological framework. Section 4 presents and discusses the empirical findings. Section 6 concludes the paper.
@@ 2 Data description and summary statistics
Throughout this research, we examine three categories of highly capitalized or actively traded assets: Firstly, AI ETFs, specifically iShares Robotics and Artificial Intelligence ETF (IRBO) and First Trust Nasdaq Artificial Intelligence and Robotics ETF (ROBT), which are representative assets in the artificial intelligence sector, with data sourced from Investing.com. Secondly, AI-based tokens, including Render (RENDER), NEAR Protocol (NEAR), Internet Computer (ICP), Filecoin (FIL), and Artificial Superintelligence Alliance (FET), obtained from CoinGecko. Among these tokens, FET was formed through the merger of Fetch.ai, SingularityNET, and Ocean Protocol, and was later rebranded as the Artificial Superintelligence token. In this study, it is referred to as FET for simplicity. These tokens represent diverse applications within the blockchain and AI ecosystems. RENDER focuses on decentralized graphics rendering. NEAR is dedicated to developing a scalable blockchain platform. ICP serves as the native token of the Internet Computer Protocol. FIL functions as the utility token for the Filecoin network, and FET is associated with an AI-powered decentralized machine learning network. Thirdly, green assets, comprising the S&P Green Bond (SPGB) and S&P Global Clean Energy Transition Index (SPGTCED), were collected from the S&P Global website. Due to data availability constraints, this study covers the period from May 25, 2021, to January 31, 2025, utilizing daily data with a total of 927 observations. All asset returns are calculated using logarithmic returns.
@@
Mean (×$\times 10^{2}$) Variance (×$\times 10^{2}$) Skewness Ex.Kurtosis JB ERS
IRBO 0.005 0.028 --0.003 1.717*** 113.810*** --12.922***
ROBT 0.002 0.024 0.069 1.264*** 62.427*** --14.157 ***
FET 0.494 0.677 1.252*** 7.151*** 2217.373*** --1.276 ***
FIL --0.054 0.474 1.738*** 17.811*** 12719.882*** --2.573***
ICP --0.054 0.504 1.352*** 10.751*** 4746.375*** --9.715 ***
NEAR 0.330 0.558 0.748*** 4.623*** 911.714*** --2.164 ***
RENDER 0.642 0.801 1.130*** 4.687*** 1045.970*** --1.498 ***
SPGB -0.019 0.023 0.247*** 1.825*** 138.021*** --6.615 ***
SPGTCED --0.061 0.025 0.454*** 2.307*** 237.196*** --14.460 ***
@@
Table 1 presents the descriptive statistics and preliminary tests for the AI and green assets. The mean returns vary significantly across assets, with RENDER and FET exhibiting the highest average returns, while FIL, ICP, SPGB, and SPGTCED display negative means. Variance suggests that AI tokens, particularly RENDER, exhibit higher volatility compared to AI ETFs and green assets. Skewness values indicate asymmetry in return distributions. While most assets exhibit positive skewness, IRBO shows slight letftward skew. Excess kurtosis values are significantly positive for all assets, implying leptokurtic distributions. The Jarque-Bera (JB) test strongly rejects normality for all assets at the 1% level. Additionally, the Elliott-Rothenberg-Stock (ERS) test results indicate strong stationarity for all series, as evidenced by significantly negative values. These findings suggest non-normal, highly volatile return distributions, particularly among AI tokens.
@@
In contrast, the correlations between SPGB (green bonds) and AI assets remain relatively weak, with values mostly below 0.3, suggesting limited interaction between these asset classes. However, SPGTCED (another green asset) shows a stronger correlation with AI assets, reaching 0.62 with ROBT and 0.58 with IRBO, suggesting some level of connectedness between sustainable investment themes and AI-driven assets. Notably, The correlation coefficients between ROBT and AI tokens, as well as those between clean energy assets and cryptos, largely fail to pass the significance test.

### v2

@@ Abstract
This paper investigates the risk spillovers among AI ETFs, AI tokens, and green markets using the $R^{2}$ decomposition method. We reveal several key insights. First, the overall transmission connectedness index (TCI) closely aligns with the contemporaneous TCI, while the lagged TCI is significantly lower. Second, AI ETFs and clean energy act as risk transmitters, whereas AI tokens and green bond function as risk receivers. Third, AI tokens are difficult to hedge and provide limited hedging ability compared to AI ETFs and green assets. Moreover, multivariate portfolios effectively reduce AI assets investment risk. Among them, the minimum correlation portfolio outperforms the minimum variance and minimum connectedness portfolios.
@@ 1 Introduction
The onset of the Fourth Industrial Revolution has ushered human society into the era of artificial intelligence (AI). This transformation has not only revolutionized daily life, work practices, and communication methods but has also profoundly impacted education, healthcare, finance, and transportation (Skilton and Hovsepian, 2018). Simultaneously, technological advancements have facilitated the evolution of financial markets, leading to the emergence of artificial intelligence related financial instruments. According to existing literature (Yousaf et al., 2024c, a, b; Jareño and Yousaf, 2023), AI-related assets constitute a class of financial instruments closely linked to artificial intelligence technologies and their broader industrial ecosystem. These assets can be broadly categorized into three groups: AI-themed exchange-traded funds (AI ETFs), publicly listed stocks of AI-related companies (AI-related stocks), and AI-integrated cryptocurrency assets (commonly referred to as AI tokens). AI ETFs allocate capital to companies in AI technology, robotics, cloud computing, and automation, offering diversified exposure to the AI sector (Yousaf et al., 2024c). AI-related stocks include companies where AI is a core business component or holds a significant role in revenue and strategic positioning, such as NVIDIA and Google (Yousaf et al., 2024a, b; Jareño and Yousaf, 2023). And AI Tokens integrate AI and blockchain technology, primarily facilitating decentralized AI computing, data sharing, and smart contract execution to support distributed AI development (Yousaf et al., 2024a, c; Jareño and Yousaf, 2023). These innovative asset classes not only represent the forefront of technological progress but have also become novel investment instruments, attracting significant interest from institutional and individual investors (Yousaf et al., 2024c).
The investment boom in the AI industry has driven the rapid expansion of computational infrastructure, providing strong momentum for technological innovation and productivity growth. However, this rapid development is heavily reliant on high-performance computing, which significantly increases energy consumption and raises environmental concerns (Jones, 2018; Vinuesa et al., 2020). At the same time, AI plays a critical role in advancing the green economy by optimizing renewable energy integration, monitoring environmental indicators, enhancing energy efficiency, and supporting sustainability initiatives (Vinuesa et al., 2020; Entezari et al., 2023; Nahar, 2024).
Amid the global push toward a low-carbon economy, financial markets are increasingly integrating environmental considerations into investment decisions and promoting green financing. As a result, green assets have gained prominence as key instruments in facilitating the transition to sustainability (Tiwari et al., 2022; Henriques and Sadorsky, 2024; Tiwari et al., 2023). Supported by favorable policies and strong growth prospects, these emerging markets are attracting rising investor interest, particularly in areas where green and AI-related technologies converge.
Given the complex relationship between artificial intelligence and green economic development, it is plausible that intricate financial interactions also exist between AI-related assets and green assets. Similarly, prior studies have documented dynamic linkages between green equities and technology stocks (Sadorsky, 2012; Henriques and Sadorsky, 2024; Tiwari et al., 2023), and green assets have been shown to offer diversification benefits (Tiwari et al., 2022; Henriques and Sadorsky, 2024). Yet little is known about whether green assets can help diversify the risks of AI-related investments.
Moreover, during periods of significant turmoil such as financial crises, the COVID-19 pandemic, and the Russia-Ukraine war, the interconnections between different financial markets tend to strengthen (Henriques and Sadorsky, 2024; Demiralay et al., 2021; Tiwari et al., 2022; Zeng et al., 2024). Despite their distinct mechanisms, these markets often experience similar impacts from such substantial external shocks. Under the current geopolitical tensions and ongoing economic uncertainties following the COVID-19 pandemic, investigating the information transmission between AI-related assets and green assets is both timely and of practical significance.
Therefore, this study systematically investigates the risk spillover effects between AI assets and green assets via the $R^{2}$ decompostion method (Balli et al., 2023; Naeem et al., 2024). Although AI-related assets encompass various forms, we specifically focus on linkages among AI ETFs, AI Tokens and green assets. AI ETFs invest in multiple companies, providing a comprehensive representation of market trends in the AI sector, and capturing investor sentiment and capital flows effectively. In contrast, AI Tokens exhibit distinct market characteristics, such as high volatility, rapid evolution, and a decentralized digital financial structure. Frequent fluctuations in trading volumes, along with frequent token listings, delistings, and mergers, contribute to the market’s dynamic nature. A recent example illustrating this volatility is Render’s rebranding from RNDR to RENDER and its migration from Ethereum to Solana. Given these unique attributes, we further assess portfolio performance of AI ETFs and AI Tokens alongside green assets. We estimate the bilateral hedging ratios and hedge effectiveness of different assets pairs. Additionally, we apply the the minimum variance portfolio (MVP) (Markowitz, 1952), and compare its performance with the minimum correlation portfolio (MCP) (Christoffersen et al., 2014), and the minimum connectedness portfolio (MCoP) (Broadstock et al., 2022).
This paper has the following contributions: First, the development of green investment markets is essential for addressing climate change and promoting sustainable growth. At the same time, artificial intelligence has the potential to accelerate the global energy transition by improving efficiency and supporting innovative low-carbon solutions. Understanding the financial relationship between these two sectors can thus provide meaningful support for achieving sustainability goals. Second, this paper contributes to the literature by investigating risk spillovers between AI tokens, AI ETFs, and green assets. It reveals how market shocks may transmit across these asset classes and identifies both contemporaneous and lagged spillover effects. This approach enhances the understanding of dynamic interdependencies in increasingly integrated sustainable and tech-driven markets. Third, the study assesses the hedging effectiveness of AI-related assets and compares their performance with green assets under various investment strategies. These findings offer practical implications for portfolio diversification and asset allocation, while also helping regulators evaluate market stability through a risk transmission framework.
@@
The remainder of this paper is structured as follows. Section 2 offers a review of the related literature. Section 3 provides a description of the data sets, along with summary statistics. Section 4 provides the methodological framework. Section 5 presents the dynamic connectedness analysis. Section 6 discusses the portfolio and hedging strategies. Section 7 concludes the paper.
2 Literature review
The interconnections among AI-related assets and and other instruments have garnered significant attention, including stocks, bonds, commodities, cryptocurrencies, and electricity (Yousaf et al., 2024c; Demiralay et al., 2021; Yousaf et al., 2024b; Abakah et al., 2023; Zeng et al., 2024; Sharma et al., 2023; Billah, 2025; Huynh et al., 2020; Ali et al., 2024; Ghaemi Asl and Roubaud, 2024; Abdullah et al., 2024). Overall, the empirical results demonstrate that the interlinkages vary across time–frequency domains and under different market conditions. Demiralay et al. (2021) find that the wavelet co-movements between AI and robotics stocks and other assets are time-scale dependent, with correlations increasing significantly during the COVID-19 pandemic. Yousaf et al. (2024a) report that the connectedness between AI assets and fossil fuel markets varies across time-frequencies and quantiles, with short-term fluctuations driving shock transmission, while long-term dynamics can shift assets between net transmitter and receiver roles. Abakah et al. (2023) examine the directional predictability between Bitcoin, fintech, and AI stocks. The results show that this predictability changes across different time lags and shows an oscillating pattern.
A growing body of research explores the interdependence among different types of AI-based assets (Yousaf et al., 2024c; Billah, 2025; Yousaf et al., 2024a; Jareño and Yousaf, 2023). Billah (2025) investigates the connections among AI-based financial assets, Sukuk, and Islamic equity indices using a Quantile VAR model. The study reveals distinct spillover patterns under extreme market conditions compared to normal periods, highlighting the uncertainty in the evolution of tail dependencies. However, the correlation between AI stocks and tokens tends to be relatively low and may vary over time, increasing during periods of economic turmoil (Jareño and Yousaf, 2023). AI ETFs consistently act as net transmitters, dominating the system across market conditions, while AI tokens serve as weak net recipients at median quantiles, with all variables alternating between net transmitters and recipients over time (Yousaf et al., 2024c). Yousaf et al. (2024a) investigate volatility interconnections between AI tokens, AI stocks, and fossil fuel markets. They report that risk transmission patterns vary across AI-related assets, highlighting their heterogeneous roles as either transmitters or recipients.
Some scholars have examined the determinants of spillovers and interactions between AI-related assets and other asset classes. Yousaf et al. (2024b) find that the impact of the factors on spillover effects varies across different market conditions. They provide evidence that assets such as gold, Bitcoin, and bonds have a significant negative impact on the connectedness between the electricity market and the artificial intelligence market. Billah (2025) examines how global factors influence asset spillovers among AI-based assets, Sukuk, and Islamic equity indices. The study finds that the impact of these factors on spillovers varies significantly depending on the prevailing market environment. Ali et al. (2024) argue that economic policy uncertainty and U.S. stock market volatility increase the interconnectedness between AI and fintech stocks, while geopolitical risk and gold market fluctuations have a negative impact on their correlations. Abdullah et al. (2024) find that global uncertainty factors significantly increase both upside and downside tail-risk connectedness among tech-industry tokens and equities.
Several studies have pointed out that AI-focused assets are emerging as promising tools for diversification, providing investors with exposure to more balanced and diversified portfolios (Jareño and Yousaf, 2023). According to Billah (2025), Sukuk and AI-based token markets are key recipients of return shock spillovers, which has important implications for portfolio diversification and risk management strategies. Yousaf et al. (2024c) find that under normal market conditions, AI tokens can serve as effective diversifiers within portfolios containing traditional assets. However, they point out that during periods of extreme market stress, neither AI tokens nor AI ETFs provide significant risk diversification benefits for other asset classes. Moreover, AI tokens have the potential to provide cost-effective hedging for traditional assets. Abdullah et al. (2024) highlight substantial opportunities for portfolio diversification when combining tech-industry tokens with tech-industry equities.
Moreover, there exits potential connections between AI- related assets and sustainable markets. The empirical findings indicate that the AI index consistently emits shocks to clean markets across all frequencies, while the S&P Global Clean Energy Index functions as a net sender of volatility spillover across all market conditions and frequencies (Zeng et al., 2024). Investors incorporating stocks of artificial intelligence and robotics companies, cryptocurrencies, and green bonds into their portfolios should be aware that certain investment risks persist (Huynh et al., 2020).
In summary, existing research has extensively explored the correlations between AI-related assets and other financial assets, with some studies further investigating the underlying factors influencing these relationships. However, there remains a lack of systematic research on the correlation and interdependence between AI assets and green assets. Furthermore, limited research has explored the role of sustainability-linked financial instruments in hedging risks and optimizing portfolio performance in relation to AI assets. Moreover, AI tokens, as a unique type of crypto asset, have introduced significant changes in global financial dynamics. Hence, our study contributes to the literature on spillovers between innovative AI-based assets and green markets. We emphasize the potential diversification benefits that AI-based assets can offer in portfolios and highlight their influence on overall portfolio performance.
@@ 3 Data description and summary statistics
Throughout this research, we examine three categories of highly capitalized or actively traded assets: Firstly, AI ETFs, specifically iShares Robotics and Artificial Intelligence ETF (IRBO) and First Trust Nasdaq Artificial Intelligence and Robotics ETF (ROBT), which are representative assets in the artificial intelligence sector, with data sourced from Investing.com.
Secondly, AI-based tokens, including Render (RENDER), NEAR Protocol (NEAR), Internet Computer (ICP), Filecoin (FIL), and Artificial Superintelligence Alliance (FET), obtained from CoinGecko. These tokens represent diverse applications within the blockchain and AI ecosystems. The selection of tokens in this study is based on market capitalization, trading volume, establishment date, and rankings as reported by CoinMarketCap. Tokens with lower rankings, insufficient market capitalization, and trading volume, or a relatively short history were excluded from consideration. Among these tokens, FET was formed through the merger of Fetch.ai, SingularityNET, and Ocean Protocol, and was later rebranded as the Artificial Superintelligence token. In this study, it is referred to as FET for simplicity. RENDER is a platform focused on decentralized graphics rendering. As of this analysis, it has transitioned from its previous name, Render (RNDR), to Render (RENDER). For consistency, we use “RENDER” as its shorthand throughout this article, with Render as the full name. NEAR is dedicated to developing a scalable blockchain platform. ICP serves as the native token of the Internet Computer Protocol. FIL functions as the utility token for the Filecoin network, and FET is associated with an AI-powered decentralized machine learning network.
Thirdly, green assets, comprising the S&P Green Bond (SPGB) and S&P Global Clean Energy Transition Index (SPGTCED), were collected from the S&P Global website. Due to the launch date of the ICP being May 10, 2021, and after excluding extreme fluctuations, the time period for this study spans from May 25, 2021, to January 31, 2025. The analysis utilizes daily data, with a total of 927 observations. All asset returns are calculated using logarithmic returns.
@@
Mean (×$\times 10^{2}$) Variance (×$\times 10^{2}$) Skewnes Ex.Kurtosis JB ERS
IRBO $0.005$ $0.028$ $-0.003$ $1.717^{***}$ $113.810^{***}$ $-12.922^{***}$
ROBT $0.002$ $0.024$ $0.069$ $1.264^{***}$ $62.427^{***}$ $-14.157^{***}$
FET $0.494$ $0.677$ $1.252^{***}$ $7.151^{***}$ $2217.373^{***}$ $-1.276^{***}$
FIL $-0.054$ $0.474$ $1.738^{***}$ $17.811^{***}$ $12719.882^{***}$ $-2.573^{***}$
ICP $-0.054$ $0.504$ $1.352^{***}$ $10.751^{***}$ $4746.375^{***}$ $-9.715^{***}$
NEAR $0.330$ $0.558$ $0.748^{***}$ $4.623^{***}$ $911.714^{***}$ $-2.164^{***}$
RENDER $0.642$ $0.801$ $1.130^{***}$ $4.687^{***}$ $1045.970^{***}$ $-1.498^{***}$
SPGB $-0.019$ $0.023$ $0.247^{***}$ $1.825^{***}$ $138.021^{***}$ $-6.615^{***}$
SPGTCED $-0.061$ $0.025$ $0.454^{***}$ $2.307^{***}$ $237.196^{***}$ $-14.460^{***}$
@@
Table 1 presents the descriptive statistics and preliminary tests for the AI and green assets. The mean returns vary significantly across assets, with FET and RENDER exhibiting the highest average returns, while FIL, ICP, SPGB, and SPGTCED display negative means. Variance suggests that AI tokens, particularly RENDER, exhibit higher volatility compared to AI ETFs and green assets. Skewness values indicate asymmetry in return distributions. While most assets exhibit positive skewness, IRBO shows slight letftward skew. Excess kurtosis values are significantly positive for all assets, implying leptokurtic distributions. The Jarque-Bera (JB) test strongly rejects normality for all assets at the 1% level. Additionally, the Elliott-Rothenberg-Stock (ERS) test results indicate strong stationarity for all series, as evidenced by significantly negative values. These findings suggest non-normal, highly volatile return distributions, particularly among AI tokens.
@@
In contrast, the correlations between SPGB (green bonds) and AI assets remain relatively weak, with values mostly below 0.3, suggesting limited interaction between these asset classes. However, SPGTCED (another green asset) shows a stronger correlation with AI ETFs, reaching 0.62 with ROBT and 0.58 with IRBO, suggesting some level of connectedness between sustainable investment themes and AI-driven assets. Notably, The correlation coefficients between ROBT and AI tokens, as well as those between clean energy assets and cryptos, largely fail to pass the significance test.